A glass prism made from a material of refractive index 1.86 has a refracting angle of 600. The prism is immersed in a water of refractive index 1.33. Determine the minimum deviation for a parallel beam of light transversing the prism


1
Expert's answer
2020-08-14T15:59:48-0400

"n=\frac{\sin(\frac{\phi+\delta}{2})}{\sin(\frac{\phi}{2})}\to"


"\delta=2\arcsin{(n\sin(\frac{\phi}{2}))}-\phi="



"=2\cdot\arcsin{(\frac{1.86}{1.33}\sin(\frac{60°}{2}))}-60°=28.72°"


The angle of minimum deviation is "28.72\u00b0".